[Brahms-l] Shift report 20040511 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue May 11 2004 - 00:04:11 EDT
------------------------------------------------------------
Date: Tue May 11 00:04:11 2004, Shift: 16:00-24:00
Supervisor: MM, with: 
Log book#15 pages 116 - 116

Shift Summary:
------------------------------------------------------------
1. Beam for 4 hours.

2. TPC and Non-TPC prdestal runs

   Two good data runs both with FS at 2.3 degrees full 
   field and B polarity.

   Run 11811  279 k trigger 2 and 190 k trigger 3
              MRS at 40 degrees and D5 at 350 B

   Run 11814  151 k trigger 2 and 152 k trigger 3
              MRS at 40 degrees and D5 at 1050 B

3. TPM1 tripped at the beginning of the store
   Magtrip goes off but magnets are holding steady
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 11 00:04:29 2004

This archive was generated by hypermail 2.1.8 : Tue May 11 2004 - 00:04:51 EDT